OpenStudy (anonymous):

how is enthalpy change occurs at constt. temperature if change in deltaH is heat absorbed or released which causes change in temperature

OpenStudy (anonymous):

@ganeshie8 @MrNood

OpenStudy (ookawaiioo):

When bonds are formed, they release energy. Heat is a form of energy

OpenStudy (anonymous):

this form of energy leads to increase in the temperature isnt it @oOKawaiiOo

OpenStudy (ookawaiioo):

If volume is expandable, then it wont lead to an increased in temperature.

OpenStudy (surry99):

I can help you but can you restate your question please?

OpenStudy (anonymous):

in my book it is written that\[\Delta U=q _{p}-p \Delta v \] at constt.temperature but all the heat is not dissipated as work unless \[\Delta U=0\] and also\[\Delta H=q _{p}\] so change in enthalpy also occurs at constt. temp. SO the question is WHY IN THE BOOK IT IS WRITTEN THAT DELTA U OCCURS AT CONSTT. TEMPERATURE ? AS ANY AMOUNT OF HEAT WOULD CHANGE THE TEMPERATURE... is the temp. not changed due to the latent heat absorbed by the substance present in the system....

OpenStudy (anonymous):

@surry99

OpenStudy (surry99):

hi, It should say constant pressure not constant temperature. By defintion: H = U + PV or U = H -PV now if pressure is constant deltaU = DeltaH - PdeltaV and it can be shown that at constant pressure deltaH = qp so at constant pressure deltaU = qp - PdeltaV